40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Wed Dec 7 21:18:35 2016, Koji, General, General, OMC placed on the table / the beam roughly aligned DSC_0082.jpgDSC_0084.jpg
    Reply  Thu Dec 8 21:17:09 2016, Koji, General, General, OMC placed on the table / the beam roughly aligned 
Message ID: 273     Entry time: Thu Dec 8 21:17:09 2016     In reply to: 272
Author: Koji 
Type: General 
Category: General 
Subject: OMC placed on the table / the beam roughly aligned 

The OMC cavity was locked. The alignment was precisely adjusted. The mode matching was optimized by the lens positions. The reflection during the lock is ~0.01 compared to the full reflection on non-resonance, meaning the mode matching is ~99%. The error signal was maximized (i.e. demod pahse was adjusted) by sweeping the modulation frequency. Note that the EOM is broad band. The modulation freq chosen today was 34.6MHz.

Some notes:

- The error signal has not been preamplified at all yet. Because of this, the reflection is very much sensitive to the input offset.

- The OMC needs wind shield to prevent from the noise caused by air turbulance.

- The laser PZT was actuated via the Thorlabs HV amp. Otherwise, the thermal path needs to be configured.

- One of the CCD monitor is dead. Needs more replacement.

- All the electronics should be moved to the rack. This required long BNC and SMA cables.

- The optical table needs cleaning.

ELOG V3.1.3-